When the length of the inextensible stnight
a simple pendulum is increased by a factor of 4, it's period is
a) increased by a factor of 4
Increased by a factor of 2
Decreased by a factor of 4
Decreased by a factor of 2
1
Expert's answer
2021-08-05T11:35:43-0400

Period of the pendulum oscillation is given by

"T=2\pi\sqrt{\frac{l}{g}}"

Hence, when the length of the inextensible stnight a simple pendulum is increased by a factor of 4, it's period is increased by a factor of 2.

Answer: (b) Increased by a factor of 2